OpenStudy (anonymous):

Find the vertex of this parabola: -x^2+6x-8

OpenStudy (anonymous):

f '(x)=-2x+6 f '(x)=0 --> 2x=6 x=3 y=f(3)=1

OpenStudy (anonymous):

are you really supposed to use calculus for this? if not your teacher will think either you lost your mind or you cheated

OpenStudy (anonymous):

if you are not taking calculus you might want to write \[-\frac{b}{2a}=-\frac{6}{2\times -1}=3\] and then use \[-3^2+6\times 3-8=-9+18-8=1\] to get the vertex is \[(3,1)\]

OpenStudy (anonymous):

It's the same method, except you're not thinking and just filling out formulas.

OpenStudy (anonymous):

like hell it is

OpenStudy (anonymous):

you do not need calculus to find the vertex of a quadratic. all you need is common sense. you write it in the form \[a(x+\frac{b}{2a})^2+k\] and make the rather trivial observation that this has a max or min when \[x=\frac{-b}{2a}\] and when it does you get k. finito

OpenStudy (anonymous):

'not thinking' is "take the derivative, set it equal zero and solve" as if i need to take a limit to do this easy problem

OpenStudy (anonymous):

Fair enough, I didn't think of that way to get -b/2a, because you also get -b/2a with the derivative. Still either way is easy.

OpenStudy (anonymous):

but of course you are right, either way you get it. ;)